### v1

# FullLoRA-AT: Efficiently Boosting the Robustness of Pretrained Vision Transformers

Abstract
In recent years, the Vision Transformer (ViT) model has gradually become mainstream in various computer vision tasks, and the robustness of the model has received increasing attention. However, existing large models tend to prioritize performance during training, potentially neglecting the robustness, which may lead to serious security concerns. In this paper, we establish a new challenge: exploring how to use a small number of additional parameters for adversarial finetuning to quickly and effectively enhance the adversarial robustness of a standardly trained model. To address this challenge, we develop the novel LNLoRA module, incorporating a learnable layer normalization before the conventional LoRA module, which helps mitigate magnitude differences in parameters between the adversarial and standard training paradigms. Furthermore, we propose the FullLoRA-AT framework by integrating the learnable LNLoRA modules into all key components of ViT-based models while keeping the pretrained model frozen, which can significantly improve the model robustness via adversarial finetuning in a parameter-efficient manner. Extensive experiments on CIFAR-10, CIFAR-100, and Imagenette demonstrate the superiority of our proposed FullLoRA-AT framework. It achieves comparable robustness with full finetuning while only requiring about 5% of the learnable parameters. This also effectively addresses concerns regarding extra model storage space and enormous training time caused by adversarial finetuning.
1 Introduction
In recent years, with the rapid development in the field of computer vision, deep learning models have made profound progress in various tasks, such as image classification [42, 17, 47], face recognition [30, 51, 11], and semantic segmentation [4, 6, 5]. Along with the increasing accessibility of computational resources and the rapid growth in the number of model parameters and training data, Vision Transformer (ViT)-based large models [13, 48, 31] have gradually become the mainstream models for various tasks [2, 65, 1].
Despite the significant success of large models, there is often more focus on performance during training, which can sometimes lead to neglecting the robustness of the model. Unfortunately, existing work has found that deep learning models are vulnerable to attacks by adversarial examples [46, 16, 26, 12]. This means that adding minimal, imperceptible perturbations to the input can easily lead to incorrect outputs from the model, whether it is CNN-based [16, 12] or ViT-based [32, 15] models.
Given this situation, we propose a new task setting: for models that have already been standardly pretrained, we aim to employ a lightweight approach to finetune them, thereby quickly enhancing their robustness against the adversarial examples. The lightweight aspect we refer to primarily concerns the number of additional model parameters that need to be trained during finetuning, as well as the additional training time required.
Many existing research works have proposed various adversarial defense methods to enhance model robustness, including adversarial training [33, 62], adversarial purification [37], adversarial example detection [27], and certifiable defenses [8], among which adversarial training is the most commonly used and effective method. Adversarial training typically involves incorporating adversarial examples into the training set, enabling the model to resist such examples during training and significantly enhancing its robustness. Traditional approaches to adversarial training typically involve either training from scratch or fully finetuning from a pretrained model. This process requires updating the entire set of model parameters to develop a robust model.
Figure 1: (a) A diagram of our proposed LNLoRA module. (b) A Vision Transformer (ViT) model utilizing our proposed FullLoRA-AT framework, which employs the LNLoRA module in the patch embedding layer, the multi-headed self-attention (MSA) block, the multi-layer perceptron (MLP) block and the MLP-based classification head. FullLoRA-AT enables the rapid enhancement of the adversarial robustness of the pretrained model through adversarial finetuning with a small number of trainable parameters.
However, when dealing with ViT-based large models, the extensive number of parameters in these models and the immense size of training datasets make it challenging to apply adversarial training due to its considerable computational overhead. Many methods, such as FreeAT [41] and FastAT [56], have adopted a single-step generation of adversarial examples in place of the multi-step generation used in traditional adversarial training, significantly reducing the time required for adversarial training. Additionally, works like AGAT [57] have addressed the high computational cost of the attention mechanism in ViT models by adopting an attention-guided partial token discarding method to enhance the efficiency of adversarial training. Although the aforementioned methods have somewhat reduced the time required for adversarial training, they still necessitate updating all parameters of the model, resulting in significant additional model storage costs, while also incurring notable losses in robustness. Furthermore, in recent years, several methods have been proposed for the lightweight finetuning of pretrained large models in downstream tasks. Parameter-efficient language model tuning methods such as Adapters [18], Prefix-tuning [28], LoRA [20], UniPELT [35], AdaLoRA [63], and Compacter [34] have been introduced for efficient finetuning of large language models (LLMs) in downstream tasks. Recently, Aurora [52] has also been proposed for efficient downstream task finetuning of large-scale multimodal foundation models. These methods effectively enhance the performance of pretrained models in downstream tasks by finetuning a small number of learnable parameters. However, there has been no work focusing on designing a parameter-efficient approach to improve the adversarial robustness of pretrained models.
To enhance the robustness of models pretrained using standard training in a lightweight manner, we draw inspiration from the LoRA method [20] and introduce the FullLoRA-AT framework. This framework aims to substantially improve the robustness of pretrained models against adversarial examples by adding only a small number of additional learnable parameters, which significantly reduces training time. Specifically, we first analyse the parameters from both adversarially and standardly trained models and reveal notable disparities in the parameters of scale and bias in the layernorm. These disparities can not be adequately addressed by the existing LoRA method. To tackle this issue, we introduce the novel LNLoRA module, which integrates a learnable layernorm layer within the existing LoRA module. Moreover, unlike the existing LoRA method which focuses only on the multi-headed self-attention (MSA) block and multi-layer perceptron (MLP) block in the Transformer layer, our approach introduces LNLoRA modules into more key components of the ViT, i.e., the patch embedding layer, MSA block, MLP block, and the classification head, as shown in Fig. 1. It is worth mentioning that our FullLoRA-AT method operates independently of existing advanced adversarial training techniques. This characteristic allows for seamless integration with others, leading to further improvement in the efficiency and performance of adversarial finetuning.
We evaluate our FullLoRA-AT method on three datasets and several ViT-based models. Extensive experiments demonstrate that our proposed FullLoRA-AT method can significantly enhance the robustness of pretrained models against adversarial examples while substantially saving on additional training parameters and accelerating the training speed. In the case of ViT-B on the CIFAR10 dataset, compared to fully finetuned models, our method demonstrates remarkable efficiency, which requires only about 5% of the learnable parameters to achieve comparable robustness, i.e., incurring a mere 1%-2% loss in accuracy under the evaluation of various attacks, e.g., CW [3], PGD [33] and AutoAttack [9].
The primary contributions of our research are summarized as follows:
1. We are the first to propose the task of rapidly enhancing the adversarial robustness of standardly pretrained models using a lightweight, parameter-efficient approach.
2. We introduce novel LNLoRA module, based on the analysis of the parameters from both adversarially and standardly trained models, to alleviate notable disparities in the parameters of scale and bias in the layernorm.
3. We propose the FullLoRA-AT method, a lightweight parameter-efficient approach to rapidly improve the robustness of the pretrained model, which significantly reduces the number of learnable parameters, thereby substantially saving on model storage space and reducing the time required for model finetuning.
2 Related Work
2.1 Vision Transformer
Vision Transformer (ViT), since its introduction by [13], has revolutionized the field of computer vision, marking a significant shift from the conventional Convolutional Neural Networks (CNNs) [17, 47, 21] that dominated the field for years. ViT employs self-attention mechanisms, typical of transformer models in natural language processing [50], to process images, which distinguishes it from the localized and hierarchical nature of CNNs [25, 42]. In image classification, ViT [13] has shown remarkable success, achieving state-of-the-art performance on benchmarks like ImageNet [40]. Later, various variants of ViT have been proposed to improve the efficiency and performance of ViT models, i.e., DeiT [48], Swin [31], MetaFormer [61] and PSLT [58]. With the great success of ViT in image classification, the application of ViT-based models has also extended to other visual tasks, such as object detection [2, 14, 53, 55], semantic segmentation [65, 22, 44, 7], image super resolution [60, 29, 64]. In the various visual tasks mentioned above, ViT-based models have achieved significant performance improvements and have gradually become mainstream models in the field of computer vision.
2.2 Adversarial Training
Existing research has proposed numerous methods for the adversarial defense of models, such as adversarial training [33], adversarial purification [37], adversarial example detection [27], and certifiable defense [8], with adversarial training being one of the most effective. Adversarial training is first introduced by PGD [33], which improves model accuracy under adversarial attacks by incorporating generated adversarial examples into the training set. Researches like TRADES [62] and MART [54] further enhance model robustness by optimizing the objective function during adversarial training. Due to the significant time required to generate adversarial examples in adversarial training, methods like FreeAT [41] and FastAT [56] have markedly accelerated this process through single-step attacks while maintaining high model robustness. For adversarial training on ViT models, recent studies [10, 43] have improved this process from the perspectives of model structure, training strategy, and training tricks. Mo et al. [36] improves the adversarial robustness of ViTs by randomly masking gradients from some attention blocks and masking perturbations on some patches during training. AGAT [57] addresses the computational expense of the attention mechanism in ViT models by discarding some tokens guided by attention, thus saving time in adversarial training. Rebuffi et al. [38] trains separate classification tokens of a ViT for the clean and adversarial domains, achieving good clean and robust accuracy, which can be further balanced through a model soup approach. AutoLoRA [59] uses the LoRA [20] mechanism to decouple features of clean and adversarial examples, thereby enhancing the robustness of pretrained large models in downstream tasks after finetuning.
Most existing adversarial training methods involve training or finetuning the entire set of model parameters. In contrast, our work freezes most of the pretrained model parameters and proposes to enhance the robustness of pretrained models in a lightweight manner through a small number of additional training parameters.
2.3 Parameter-efficient Model Finetune
Large models pretrained on extensive datasets exhibit commendable generalizability, making finetuning on such models for downstream tasks a prevalent paradigm. In recent years, numerous parameter-efficient finetuning methods have been proposed for efficiently optimizing models for downstream tasks. In natural language processing (NLP), methods such as Adapters [18], Prefix-tuning [28], LoRA [20], UniPELT [35], AdaLoRA [63] and Compacter [34] have been introduced. They add a small number of additional learnable parameters to the pretrained model to facilitate rapid and efficient finetuning for downstream NLP tasks. In visual tasks, Visual Prompt [23] introduces only a small number of trainable parameters in the input space while keeping the model backbone frozen, often outperforming full finetuning. Furthermore, with the recent rise of multimodal models, many works have focused on the parameter-efficient finetuning of these models. VL-Adapter [45] introduces adapter-based parameter-efficient transfer learning techniques to vision-language models, and Aurora [52] applies CP decomposition to model parameters and proposes the informative context enhancement and gated query transformation module for lightweight finetuning.
Existing parameter-efficient model finetuning methods are based on pretrained models and target specific downstream tasks. Our work, however, is the first to propose using a parameter-efficient approach to enhance the robustness of standardly trained models against adversarial examples.
3 Method
In this section, we first provide a brief overview of the Vision Transformer [13] (ViT) model, which has been widely used in various computer vision tasks, as well as the existing Low-Rank Adaptation (LoRA) method [20], known for its efficiency in lightweight finetuning of pretrained models for various downstream tasks. Following this, we introduce our novel FullLoRA-AT method, a lightweight approach designed to enhance the adversarial robustness of standardly pretrained models in a parameter-efficient manner.
3.1 Preliminary
This section begins with a brief overview of the Vision Transformer (ViT) model structure and the existing Low-Rank Adaptation (LoRA) method.
3.1.1 Vision Transformer
In recent years, the Vision Transformer (ViT) has achieved remarkable performance improvements in various visual tasks and has gradually become the mainstream model for these tasks.
The ViT model first resizes the input image $x\in\mathbb{R}^{H\times W\times C}$ into different patches $x_{p}\in\mathbb{R}^{N\times(P^{2}\cdot C)}$, where $(H,W)$ is the resolution of the image, $C$ is the number of channels in the image, $(P,P)$ is the resolution of each image patch, and $N=HW/P^{2}$ is the number of patches after resizing. ViT initially maps these image patches to features through a patch embedding layer $E\in\mathbb{R}^{(P^{2}\cdot C)\times D}$ , concatenates them with a special classification token $x_{class}$, and adds position embedding $E_{pos}\in\mathbb{R}^{(N+1)\times D}$ to obtain the embedded patches $z_{0}$. These embedded patches are then transformed by $L$ consecutive Transformer layers, each consisting of a multi-headed self-attention (MSA) and multi-layer perceptron (MLP) block, resulting in the output $z_{L}$ of the entire Transformer encoder. Finally, the feature $z_{L}^{0}$ at the classification token position of $z_{L}$ is processed through an MLP-based classification head layer to produce the model’s final classification result $y$. The overall process can be formulated as follows:
$z_{0}=[x_{\text{class}};x_{p}^{1}E;x_{p}^{2}E;\cdots;x_{p}^{N}E]+E_{pos},$ (1)
$z_{l}^{\prime}=\mathrm{MSA}(\mathrm{LN}(z_{l-1}))+z_{l-1},\quad\mathrm{~{}for~{}}l=1,\ldots,L$ (2)
$z_{l}=\mathrm{MLP}(\mathrm{LN}(z_{l}^{\prime}))+z_{l}^{\prime},\quad\mathrm{for~{}}l=1,\ldots,L$ (3)
$y=\mathrm{MLPHead}(z_{L}^{0}),$ (4)
where $\mathrm{LN}$ is the layernorm, $\mathrm{MSA}$ is the multi-headed self-attention block, $\mathrm{MLP}$ is the multi-layer perceptron block and $\mathrm{MLPHead}$ is the MLP-based classification head.
3.1.2 Low-Rank Adaptation
Low-rank adaptation (LoRA) is an efficient method for finetuning pretrained models on downstream tasks. It utilizes a low-rank approach for updating additional model weights rather than directly finetuning the pretrained model, effectively reducing the number of parameters and training costs.
For a pretrained weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA constrains the update during finetuning for downstream tasks by representing the parameter update with a low-rank decomposition: $W+\Delta W=W+BA$, where $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$, and the rank $r\leq min(d,k)$. During the finetuning, the pretrained weight matrix $W$ remains frozen, while $A$ and $B$ serve as the learnable parameters. The LoRA module is typically used in the MSA block and the MLP block of Transformer layers to facilitate parameter-efficient finetuning of pretrained models.
(a) $\|W_{clean}\|$
(b) $\|W_{adversarial}\|$
(c) $\|W_{clean}-W_{adversarial}\|$
Figure 2: Visualization of the average absolute values of different parameters across the 12 Transformer layers of the ViT model. (a) presents the model with standard training, (b) presents the model with adversarial training, and (c) presents the differences in model parameters between (a) and (b).
3.2 FullLoRA Adversarial Training
Existing Vision Transformer (ViT) models have achieved promising results in various visual tasks. However, during model training, there is often a greater focus on performance, potentially leading to a neglect of the model’s robustness. In response to this situation, we propose a new task setting: for pretrained models obtained through standard training, we aim to enhance the adversarial robustness quickly using a parameter-efficient approach for finetuning. To this end, we introduce FullLoRA Adversarial Training, a method that minimally increases parameters to enhance model robustness. Compared to the commonly used full finetuning approach, our FullLoRA-AT introduces a significantly smaller number of additional training parameters, thereby saving on both training time and model storage costs during finetuning.
In the following, we will provide a detailed introduction to our proposed FullLoRA-AT method.
3.2.1 LNLoRA
We first analyze the differences in parameters between models obtained by standard training and adversarial training. We conduct the experiment on the CIFAR10 dataset [24] with the ViT-B model [13]. As shown in Fig. 2, comparing the average absolute values of different parameters across the 12 Transformer layers, we can clearly see that the scale and bias coefficients in the layernorm preceding each MSA and MLP block in the ViT exhibit particularly significant differences between the two models, especially can be observed from (c) in Fig. 2, which represents their residuals.
Due to the feature scaling functionality of the layernorm, the features of clean images and adversarial examples in models with standard and adversarial training show significant differences after scaling through the layernorm before entering the MSA and MLP blocks. The existing LoRA method [20] applies only to the weights in the Key, Query, and Value parameters of the MSA block and the two fully connected layers of the MLP block. When freezing the parameters of the standard training model and only applying the existing LoRA method for adversarial finetuning, it becomes challenging for the model to approximate the parameters obtained through fully adversarial training using the low-rank matrices in the LoRA module.
To address these limitations, we introduce the LNLoRA module, which incorporates an additional layernorm preceding the existing LoRA module. The purpose of LNLoRA is to learn the differences in feature magnitude between standard and adversarial trained models via the extra layernorm, allowing the LNLoRA module to better approximate the differences between the parameters of the two models. Consider an image feature represented as $x$, undergoing an initial matrix transformation $h=Wx$. With the addition of our proposed layernorm-included LNLoRA module, it yields:
$h=Wx+BA\cdot\textrm{LN}(x)=Wx+BA(\alpha x+\beta),$ (5)
where $\alpha$ and $\beta$ denote the scale and bias coefficients in the layernorm. The integration of layernorm into the LNLoRA module significantly enhances the model’s ability to adapt to adversarial examples, thereby boosting its robustness with a small number of learnable parameters efficiently.
Table 1: Ablation study on the use of LoRA/LNLoRA in different components of the ViT model.
Method Setting Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack Time
Patch Embedding MSA block MLP block Classification Head LoRA/LNLoRA Paras (M) (%) (%) (%) (%) (%) (h)
Fully Finetune 85.15 82.83 50.26 52.85 52.56 48.26 11.17
LoRA [20] $\surd$ $\surd$ LoRA 4.72 84.11 47.14 49.96 49.74 45.13 8.10
(a) $\surd$ $\surd$ $\surd$ LoRA 4.74 84.36 47.53 50.31 50.20 45.59 8.17
(b) $\surd$ $\surd$ $\surd$ LoRA 4.74 84.21 47.68 50.49 50.31 45.67 8.18
(c) $\surd$ $\surd$ $\surd$ $\surd$ LoRA 4.77 85.55 48.75 50.83 50.43 46.26 8.23
FullLoRA-AT (ours) $\surd$ $\surd$ $\surd$ $\surd$ LNLoRA 4.85 87.12 49.76 51.38 50.83 46.94 8.38
3.2.2 FullLoRA-AT
Based on the aforementioned LNLoRA module, we introduce the FullLoRA-AT method, a novel and lightweight approach designed to enhance the adversarial robustness of standardly pretrained models in a parameter-efficient manner. As depicted in Fig. 1, the LNLoRA module is seamlessly integrated into several key components of the ViT model, including the patch embedding layer, MSA block, MLP block, and the classification head.
Traditional LoRA methods predominantly focus on modifying parameters within the MSA and MLP blocks of the ViT model, while neglecting the parameters of other model components. However, given the significant differences in image features between clean and adversarial examples, it becomes essential to adaptively finetune the low-level layers of the model, which may facilitate the model’s ability to effectively learn features from adversarial examples that are crucial for accurate classification. Therefore, we extend the LNLoRA module to also encompass the patch embedding layer of the ViT model.
Additionally, given the critical role of parameters in the classification head layer for determining classification outcomes, and the usual practice of unfreezing this layer during finetuning, we also apply the LNLoRA module to the classifier head layer. This extension aims to facilitate a more parameter-efficient adaptation of the model when confronting adversarial examples. Our subsequent ablation studies in the experiment further reveal that the integration of LNLoRA across these key components of ViT model, as opposed to the traditional LoRA method, clearly further boosts the robustness of the model against adversarial attacks.
It is worth mentioning that our FullLoRA-AT method operates independently of existing advanced adversarial training techniques. This characteristic allows for seamless integration with others, leading to further improvement in the efficiency and performance of adversarial finetuning.
4 Experiment
In this section, we first introduce the experimental setting, including datasets, models, comparison methods, and evaluation metrics. Subsequently, we conduct the ablation study to explore the impact of employing LoRA and LNLoRA modules on different components of the ViT model on model robustness. This is followed by comparisons with existing methods on various models and datasets. We then analyze the impact of using different rank $r$ values on the effectiveness of FullLoRA adversarial training. Finally, we integrate FullLoRA-AT with other advanced methods to further accelerate the adversarial training and improve the robustness of the model, effectively demonstrating the generalization of our approach.
4.1 Experimental Setting
Dataset. We conduct experiments on three commonly used datasets in the field of adversarial training, i.e., CIFAR10 [24], CIFAR100 [24] and Imagenette [19]. CIFAR10 and CIFAR100 each contains images of 10 and 100 classes respectively, with training and test datasets comprising 50,000 and 10,000 images respectively. The image size in these datasets is 32$\times$32. Imagenette, a subset of ImageNet [40], includes 10 classes of images, with training and test datasets consisting of 13,000 and 500 images respectively. The image size in Imagenette is 224$\times$244.
Models. We choose ViT-S [13], ViT-B [13] and Swin-B [31] models to conduct the experiments.
Comparison methods. In our study, we compare our FullLoRA-AT method with other approaches such as LoRA [20], UniPELT [35], and Aurora [52], which are all methods for parameter-efficient finetuning on pretrained models. Additionally, following [38], we also compare our approach against strategies that exclusively update specific components during finetuning, e.g., the patch embedding layer and the scale and offset parameters of all layernorms.
Evaluation metrics. In our evaluation of the robustness of various models against adversarial examples, we employ accuracy metrics under different adversarial attack methods. Specifically, we utilize CW-20 [3], PGD-20, PGD-100 [33], and AutoAttack [9] to evaluate model robustness. CW-k and PGD-k denote CW and PGD attacks with k iterative steps, respectively. The perturbation magnitude for these attacks is set to 8/255, with a step size of 1/255. Additionally, we evaluate the efficiency of finetuning pretrained models for robustness by considering the number of learnable parameters required during finetuning and the time taken for the finetuning process.
Details. In our experiments, finetuning on all datasets is conducted over 40 epochs, with the learning rate being scaled down by a factor of 0.1 after the 35-th and 38-th epochs. In default, we employ standard adversarial training with a PGD-10 attack as our baseline method. The SGD optimizer [39] is utilized for the training process. The learning rate settings vary depending on the finetuning method employed: for fully finetuned models and Aurora [52], a learning rate of 0.1 is used; for LoRA [20] and our FullLoRA-AT, the learning rate is set to 5; for UniPELT [35], the learning rate is 0.5; and for the remaining methods, a learning rate of 1 is applied. In the case of LoRA and FullLoRA-AT, considering the trade-off between model performance and training efficiency, we set the rank $r$ to 32 by default. The experiments are carried out on a single NVIDIA GeForce RTX 3090. For CIFAR10 and CIFAR100 datasets, the batch size is set to 512, and for Imagenette, the batch size is 64.
4.2 Ablation Study
We conduct the ablation study using LoRA and LNLoRA modules on different components of the ViT model, and the results are presented in Tab. 1. The experiment is conducted using the ViT-B model on the CIFAR10 dataset. As clearly shown, compared to the original LoRA method which only adds low-rank matrices in the MSA block and MLP block for finetuning, the application of the LoRA module additionally in the classification head and patch embedding layer respectively (i.e., (a) and (b) in Tab. 1) improves the robustness of model under various attacks. Employing the LoRA module across all four mentioned components (i.e., (c) in Tab. 1) further enhances the robustness of the model. Moreover, replacing the LoRA module with our proposed LNLoRA module can further improve the robustness, demonstrating the greater suitability of our newly proposed LNLoRA module for adversarial finetuning tasks on pretrained models. Additionally, as different modules are incrementally added, not only does the robustness of the model improve, but the clean accuracy of the model for classifying clean images also increases consistently.
4.3 Comparison with Other Methods
We first compare the robustness against different attack methods of pretrained models finetuned using our proposed FullLoRA-AT method with other lightweight finetuning methods on various datasets. The results are shown in Tab. 2. When only the patch embedding layer or the scale and offset parameters of layer normalization are used for adversarial finetuning, the number of trainable parameters is minimal and the training speed is relatively fast, whereas the robustness of the model in this scenario is weak. While Aurora [52], LoRA [20], and UniPELT [35] methods show improvements in model robustness, the Aurora method involves a large number of additional matrix multiplication operations due to CP decomposition of the weight matrices, leading to a longer training time, even exceeding that of full finetuning. Also, the UniPELT method adds a considerable amount of extra parameters, resulting in unsatisfactory training duration.
Our proposed FullLoRA-AT method, at rank $r=32$, our FullLoRA-AT method shows a robustness improvement of 2% $\sim$ 4% on average across different datasets compared to the LoRA method, with similar parameter count and training duration, which clearly demonstrates the effectiveness of our approach. When the rank is decreased to $r=16$, our FullLoRA-AT can still achieves similar or even better model robustness compared to the LoRA method while requiring only half the number of parameters for finetuning and reducing training time. Moreover, compared to fully finetuning the entire model, our method uses only about 5% of the parameters and 75% of the training duration to achieve comparable model robustness. On the dataset of CIFAR-10, our method incurs an average loss of only about 1% in robust accuracy, yet it simultaneously achieves a 4.29% improvement in clean accuracy.
We also conduct experiments on different ViT-based models using the CIFAR10 dataset, and the results are displayed in Tab. 3. Similarly, our method achieves the best results across different models, demonstrating the superiority and generalization of our proposed FullLoRA-AT method.
Table 2: Comparison with other methods across different datasets.
Model Method Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack Time
Paras (M) (%) (%) (%) (%) (%) (h)
CIFAR-10 Fully Finetune 85.15 82.83 50.26 52.85 52.56 48.26 11.17
Patch Embedding Layer 0.04 41.14 15.77 19.91 17.72 15.13 7.42
Layer Normalization 0.04 35.64 20.87 23.58 23.57 20.97 7.79
Aurora [52] 0.09 76.09 38.50 41.85 41.67 39.87 11.48
LoRA (r=32) [20] 4.72 84.11 47.14 49.96 49.74 45.13 8.10
UniPELT [35] 11.42 84.63 47.63 49.49 48.99 44.63 13.11
FullLoRA-AT (ours, r=16) 2.46 85.93 47.68 50.33 49.46 45.82 8.01
FullLoRA-AT (ours, r=32) 4.85 87.12 49.76 51.38 50.83 46.94 8.38
CIFAR-100 Fully Finetune 85.22 61.02 28.61 30.58 30.43 27.49 11.18
Patch Embedding Layer 0.04 26.31 10.67 11.98 11.81 9.51 7.42
Layer Normalization 0.04 34.37 13.16 13.79 13.72 10.84 7.80
Aurora [52] 0.09 55.17 19.27 19.81 19.73 17.52 11.48
LoRA (r=32) [20] 4.72 63.91 24.58 25.83 25.71 22.53 8.11
UniPELT [35] 11.42 55.38 21.63 24.69 24.55 22.07 13.11
FullLoRA-AT (ours, r=16) 2.46 62.34 24.99 26.31 26.19 23.69 8.00
FullLoRA-AT (ours, r=32) 4.85 61.20 25.86 28.40 28.23 25.78 8.38
Imagenette Fully Finetune 85.15 93.40 69.00 68.80 68.60 67.60 8.52
Patch Embedding Layer 0.04 37.60 23.40 25.60 25.60 24.70 5.76
Layer Normalization 0.04 61.40 23.60 28.80 28.80 27.60 5.89
Aurora [52] 0.09 87.10 51.30 53.20 52.90 51.80 9.07
LoRA (r=32) [20] 4.72 94.00 63.80 62.00 61.20 60.00 6.15
UniPELT [35] 11.42 91.40 58.00 57.80 56.20 55.30 9.58
FullLoRA-AT (ours, r=16) 2.46 93.70 64.10 63.70 63.00 62.10 6.01
FullLoRA-AT (ours, r=32) 4.85 93.40 66.20 66.20 65.60 64.00 6.32
Table 3: Comparison with other methods on different ViT-based models.
Model Method Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack Time
Paras (M) (%) (%) (%) (%) (%) (h)
ViT-B Fully Finetune 85.15 82.83 50.26 52.85 52.56 48.26 11.17
Aurora [52] 0.09 76.09 38.50 41.85 41.67 39.87 11.48
LoRA (r=32) [20] 4.72 84.11 47.14 49.96 49.74 45.13 8.10
UniPELT [35] 11.42 84.63 47.63 49.49 48.99 44.63 13.11
FullLoRA-AT (ours, r=16) 2.46 85.93 47.68 50.33 49.46 45.82 8.01
FullLoRA-AT (ours, r=32) 4.85 87.12 49.76 51.38 50.83 46.94 8.38
ViT-S Fully Finetune 47.34 79.94 48.02 50.84 50.62 46.22 5.31
Aurora [52] 0.08 70.52 36.12 38.91 38.79 35.24 5.46
LoRA (r=32) [20] 3.15 78.31 44.87 46.10 45.98 41.37 3.87
UniPELT [35] 7.75 73.40 39.82 43.68 43.49 40.09 5.93
FullLoRA-AT (ours, r=16) 1.65 78.66 45.10 46.93 46.60 42.31 3.69
FullLoRA-AT (ours, r=32) 3.25 79.03 45.63 48.58 48.42 44.93 3.92
Swin-B Fully Finetune 86.69 81.29 48.36 50.99 50.88 46.57 6.44
Aurora [52] 0.09 75.13 37.52 38.17 38.03 35.26 4.79
LoRA (r=32) [20] 6.16 80.94 45.32 46.50 46.37 43.08 4.66
UniPELT [35] 11.58 79.55 44.16 45.28 44.93 41.39 5.07
FullLoRA-AT (ours, r=16) 3.17 82.86 46.10 47.53 47.34 43.57 4.49
FullLoRA-AT (ours, r=32) 6.29 83.67 46.73 48.16 48.05 44.16 4.81
Table 4: Results of our FullLoRA-AT under different rank $r$.
Method Rank $r$ Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack Time
Paras (M) (%) (%) (%) (%) (%) (h)
Fully Finetune 85.15 82.83 50.26 52.85 52.56 48.26 11.17
FullLoRA-AT (ours) 8 1.27 84.52 46.60 49.26 48.91 45.10 7.87
16 2.46 85.93 47.68 50.33 49.46 45.82 8.01
32 4.85 87.12 49.76 51.38 50.83 46.94 8.38
64 9.62 86.82 50.25 51.50 51.19 47.12 8.72
4.4 Results under different rank $r$
We compare the robustness of pretrained models adversarially finetuned using our proposed FullLoRA-AT method with different ranks $r$ in the LNLoRA module. As shown in Tab. 4, with the increase of $r$, there is a corresponding increase in the number of trainable parameters during finetuning, as well as a slight increase in the required training duration. Concurrently, the robustness of the resulting model also noticeably improves. When $r=8$, the model finetuned using our FullLoRA-AT method, with only about 1.5% of the trainable parameters used for full finetuning, incurs only a 3%-4% loss in robustness under various attacks. Increasing $r$ to 64, which utilizes 11.3% of the training parameters, further reduces the robustness loss to approximately 1%. Considering the trade-off between the number of parameters used for adversarial finetuning and the robustness of the model, we choose rank $r=32$ as the default for other experiments in our work.
4.5 Combined with other advanced adversarial training methods
Our method can also be further combined with other advanced adversarial training methods to enhance the results of our adversarial finetuning.
In Tab. 5, we present the results of combining FullLoRA-AT with methods designed to accelerate the speed of adversarial training. In our previous experiments, we use a 10-step PGD [33] to generate adversarial examples for training, referred to as PGD-10. When we switch to PGD-5, the time required for adversarial finetuning on pretrained models is reduced, and our method consistently outperforms the LoRA [20] in terms of robustness. Similarly, when combined with the AGAT [57], an attention-guided tokens discarding approach, the adversarial finetuning time is further reduced, albeit at the expense of a slight decrease in model robustness. In summary, when combined with different advanced methods for accelerating adversarial training, our FullLoRA-AT consistently achieves higher model robustness on various datasets compared to the LoRA method, demonstrating its broad applicability.
Additionally, we also combine our FullLoRA-AT with other methods aimed at enhancing the performance of adversarial robustness. As shown in Tab. 6, when combined with MART [54] and PRM [36], our FullLoRA-AT method still significantly outperforms the LoRA method in terms of model robustness across different datasets.
Table 5: Results of combining FullLoRA-AT with other speed-accelerating adversarial training methods.
Dataset Combined Method Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack Time
Method Paras (M) (%) (%) (%) (%) (%) (h)
CIFAR-10 PGD-5 [33] Fully Finetune 85.15 86.98 49.97 51.08 50.52 47.58 6.37
LoRA [20] 4.72 86.10 45.09 45.23 45.01 42.46 4.10
FullLoRA-AT (ours) 4.85 89.30 47.66 47.92 47.18 44.97 4.36
AGAT [57] Fully Finetune 85.15 87.42 45.02 47.75 47.54 42.86 5.76
LoRA [20] 4.72 86.34 42.68 43.89 43.65 39.68 3.81
FullLoRA-AT (ours) 4.85 89.47 44.07 45.00 44.45 41.03 3.93
CIFAR-100 PGD-5 [33] Fully Finetune 85.22 63.27 27.21 27.64 27.57 25.88 6.39
LoRA [20] 4.72 65.16 23.49 24.87 24.63 21.06 4.11
FullLoRA-AT (ours) 4.85 66.49 24.75 26.10 26.01 24.49 4.38
AGAT [57] Fully Finetune 85.22 63.91 25.36 25.74 25.61 24.37 5.77
LoRA [20] 4.72 65.18 21.73 21.99 21.86 20.42 3.82
FullLoRA-AT (ours) 4.85 66.31 22.93 23.07 23.01 22.54 3.93
Imagenette PGD-5 [33] Fully Finetune 85.15 94.40 64.60 64.20 63.60 62.80 4.46
LoRA [20] 4.72 93.10 60.70 58.90 58.30 57.60 3.31
FullLoRA-AT (ours) 4.85 94.00 63.80 62.00 61.20 60.30 3.37
AGAT [57] Fully Finetune 85.15 94.70 61.90 61.60 61.10 60.60 4.15
LoRA [20] 4.72 93.00 57.20 56.20 56.00 55.10 3.19
FullLoRA-AT (ours) 4.85 95.20 60.20 60.00 59.70 58.90 3.23
Table 6: Results of combining FullLoRA-AT with other performance-enhancing adversarial training methods.
Dataset Combined Method Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack Time
Method Paras (M) (%) (%) (%) (%) (%) (h)
CIFAR-10 MART [54] Fully Finetune 85.15 83.65 51.09 53.62 53.41 49.02 12.13
LoRA [20] 4.72 84.82 47.93 50.77 50.55 46.01 8.86
FullLoRA-AT (ours) 4.85 87.99 50.59 52.09 51.64 47.66 8.97
PRM [36] Fully Finetune 85.15 83.43 50.97 53.49 53.12 48.81 11.23
LoRA [20] 4.72 84.62 47.71 50.53 50.45 45.78 8.16
FullLoRA-AT (ours) 4.85 87.76 50.39 51.82 51.42 47.36 8.39
CIFAR-100 MART [54] Fully Finetune 85.22 61.82 29.51 31.46 31.20 28.23 12.15
LoRA [20] 4.72 61.70 25.44 26.59 26.60 23.36 8.88
FullLoRA-AT (ours) 4.85 62.00 26.60 29.22 29.13 26.68 8.99
PRM [36] Fully Finetune 85.22 61.46 29.15 31.12 31.12 28.13 11.25
LoRA [20] 4.72 62.59 25.11 26.26 26.17 22.97 8.18
FullLoRA-AT (ours) 4.85 61.81 26.35 28.88 28.70 26.35 8.40
Imagenette MART [54] Fully Finetune 85.15 94.26 69.81 69.68 69.40 68.49 9.32
LoRA [20] 4.72 94.06 64.57 62.78 62.00 60.88 6.75
FullLoRA-AT (ours) 4.85 94.19 66.99 67.10 66.51 64.91 6.83
PRM [36] Fully Finetune 85.15 93.95 69.68 69.47 69.07 68.22 8.59
LoRA [20] 4.72 94.07 64.48 62.54 61.63 60.49 6.19
FullLoRA-AT (ours) 4.85 94.04 66.69 66.83 66.22 64.64 6.35
5 Conclusion
In conclusion, to achieve parameter-efficient robustness finetuning for standard pretrained models, our work presents the FullLoRA-AT method, an innovative and lightweight approach designed to significantly enhance the adversarial robustness of pretrained ViT models in a parameter-efficient manner. By integrating the newly proposed LNLoRA modules into several key components of the ViT model, our method can rapidly improve the robustness of the pretrained model with a small number of extra parameters. Extensive experiments show that our method not only addresses critical limitations of previous approaches but also showcases remarkable improvement in the robustness of models against adversarial attacks.
In the future, we plan to apply our method to larger-scale models, such as LLaMA [49], to further validate its effectiveness in more realistic scenarios.